# Loss Shaping Constraints for Long-Term Time Series Forecasting

### v1

@@ 1 Introduction
Predicting multiple future values of time series data, also known as multi-step forecasting (Bontempi et al., 2013), has a myriad of applications such as weather (Wang et al., 2016), demand (Yi et al., 2022), price (Chen et al., 2018) and movility (Bai et al., 2019) forecasting. Several approaches to generating predictions for the next window have been proposed, including direct (Chevillon, 2007; Sorjamaa et al., 2007), autoregressive or recursive (Hill et al., 1996; Tiao & Tsay, 1994; Hamzaebi et al., 2009) and MIMO (Kline, 2004; Bontempi, 2008; Bontempi & Ben Taieb, 2011; Kitaev et al., 2019; Zhou et al., 2021; Wu et al., 2021; Liu et al., 2021; Zhou et al., 2022; Nie et al., 2022; Zeng et al., 2023; Garza & Mergenthaler-Canseco, 2023; Das et al., 2023b) techniques. Moreover, a plethora of learning parametrizations exist, ranging from linear models (Box & Jenkins, 1976; Zeng et al., 2023) to recent transformer architectures (Kitaev et al., 2019; Zhou et al., 2021; Wu et al., 2021; Liu et al., 2021; Zhou et al., 2022; Nie et al., 2022; Garza & Mergenthaler-Canseco, 2023; Das et al., 2023b).
@@
Recent works using transformer based architectures focus on aggregate metrics, see for example (Kitaev et al., 2019; Wu et al., 2021; Nie et al., 2022), while addressing errors at different time steps has recieved little attention (Cheng et al., 2023). Our first contribution is providing an empirical analysis in this direction:
  • (C1)
    We provide an empirical analysis of the distribution of errors across the prediction windows using popular models (Kitaev et al., 2019; Zhou et al., 2021; Wu et al., 2021) as well as a vanilla transformer in common benchmark datasets.
The analysis in (C1) gives evidence supporting that loss landscapes vary considerably between datasets and architectures. However, having no control on the error incurred at each time step in the predicted window can be detrimental in scenarios that can be affected by this variability. For example, analysing average behaviour is insufficient to assess financial risks in econometrics  (Chavleishvili & Manganelli, 2019) or to ensure stability when using the predictor in a Model Predictive Control framework (Terzi et al., 2021).
Therefore, our second contribution is addressing this problem through constrained learning.
@@
Since this constrains the distribution of the loss per time step across the window, we call this approach loss shaping constraints. Due to the challenges in finding appropriate constraints, we leverage a recent approach (Hounie et al., 2023) to re-interpret the specification as a soft constraint and find an optimal relaxation while jointly solving the learning task. We leverage recent duality results (Chamon & Ribeiro, 2020) to show that, despite being non-convex, the resulting problem has a bounded duality gap. This allows us to provide approximation guarantees at the step level.
@@ 1.1 Related work
In the context of transformer architectures for time series forecasting, to the best of our knowledge, loss at the step level has only been reported in (Cheng et al., 2023). Since the focus of the analysis is portraying the effect of a multi task learning approach on loss imbalance, the number of models and datasets for which step-wise errors across the window are reported is limited. Thus, it does not allow comparisons across datasets and models, and does not uncover the erratic (and non monotonic) behaviour that we have found empirically for some datasets and models. Recent works in generative time series models have also sought to impose constraints using penalty based methods (Coletta et al., 2023), but the nature of the constraints and proposed approach also differs from ours.
@@ 2 Multi-Step Time Series Forecasting
$~{}\operatorname{min.}_{\boldsymbol{\theta}\in\Theta}\quad\mathbb{E}_{(\mathbf{x}_{c},y_{p})\sim\mathfrak{D}}\left[\frac{1}{T_{p}}\sum_{i=1}^{T_{p}}\ell_{i}({f_{\boldsymbol{\theta}}(\mathbf{x}_{c})},y_{p})\right],$ (ERM)
@@
However, this choice of objective does not account for the structure or distribution of the errors across different time steps, which can lead to disparate behaviour across the predicted window as depicted in Figure 1. In particular, we observe empirically that state-of-the-art transformer architectures (Kitaev et al., 2019; Zhou et al., 2021; Wu et al., 2021) can yield highly varying loss dynamics, including non-monotonic, flat and highly non-linear landscapes, as presented in Section 5.1.
@@ 3 Loss Shaping Constraints
3.1 Adapting Constraints: A Resilient Approach.
An important challenge is the specification of loss shaping constraints that are not overly restrictive and make the problem feasible. Whereas in unconstrained average risk minimization (ERM) an optimal function $\theta^{\star}$ always exists, in (3) there may be no parameters in $\Theta$ satisfy the requirements. In practice, landing on satisfiable loss shaping requirements may necessitate the relaxation of some constraints –i.e., the values of $\boldsymbol{\epsilon}_{i}$ in (3)–. This is challenging because assessing the impact of tightening or relaxing a particular constraint can have intricate dependencies with the model class, the unknown data distribution, and learning algorithm, which are hard to determine a priori. Thus, the point at issue is coming up with reasonable constraints that achieve a desirable trade-off between controlling the distribution of the error across the predictive window and attaining good average performance. That is, for larger constraint levels the average performance improves, although the constraints will have less impact on the optimal function $f_{\theta^{\star}}$.
In order to do so, we introduce a non negative perturbation $\boldsymbol{\zeta}_{t}\in\mathbb{R}_{+}$ associated to each time step, and consider relaxing the $i-th$ constraint in the original problem by $\boldsymbol{\zeta}_{i}$. Explicitly, we impose $\mathbb{E}_{(\mathbf{x}_{c},y_{p})\sim\mathfrak{D}}\left[\ell_{i}({f_{\boldsymbol{\theta}}(\mathbf{x}_{c})},y_{p})\right]\leq\boldsymbol{\epsilon}_{i}+\boldsymbol{\zeta}_{i}$. We also introduce a differentiable, convex, non-decreasing cost $h:\mathbb{R}_{+}^{T_{p}}\to\mathbb{R}_{+}$, that penalizes deviating from the original specification, e.g., the squared $L_{2}$ norm $h(\boldsymbol{\zeta})\propto\|\boldsymbol{\zeta}\|_{2}^{2}$.
This allow us to re-interpret the initial constraint levels $\boldsymbol{\epsilon}$ as a soft constraint and use the perturbation cost $h$ to search for a perturbation $\zeta$ that achieves a good trade-off with average performance. Therefore, we propose to solve a Resilient Constrained Learning problem (Hounie et al., 2023), which is defined as follows:
[…]
The key property of the optimal relaxation $\boldsymbol{\zeta}^{\star}$ is that it equates the marginal cost of the relaxation (as measured by $h$) with the marginal cost of its effect on the optimal cost (Hounie et al., 2023). The rationale behind introducing a learnt relaxation of constraint levels is making the problem easier to solve. In the next section, we introduce practical primal-dual algorithms that enable approximating the constrained (3) and resilient (3.1) problems based on samples and finite (possibly non-convex) parametrizations such as transformer architectures.
@@ 4 Empirical Dual Resilient and Constrained Learning
A challenge in solving (3) and (3.1) is that in general, (i) no closed form projection into the feasible set or proximal operator exists, and (ii) it involves an unknown data distribution $\mathfrak{D}$. In what follows, we describe our approach for tackling the Resilient problem, noting that Primal-Dual algorithms tackling the original constrained problem can be similarly derived by simply excluding the slack variables.
To undertake 3.1, (i) we replace expectations by sample means over a data set $\{(\boldsymbol{x}_{c}^{n},y_{p}^{n})\>:\>n=1,\cdots,N\}$, as typically done in (unconstrained) statistical learning, and (ii) resort to its Lagrangian dual.
@@
where $\hat{L}$ is the empirical Lagrangian of (3.1), defined as
${\hfilneg\hat{L}(\boldsymbol{\boldsymbol{\theta}},\boldsymbol{\lambda},\boldsymbol{\epsilon},\boldsymbol{\zeta}):=\hskip 10000.0pt}$
@@
The Empirical Dual problem (ED-LS) is an approximation of the Dual problem associated to 3.1 based on training samples. The dual problem itself can be interpreted as finding the tightest lower bound on the primal. Although the estimation of expectations using sample means and non-convexity of the hypothesis class can introduce a duality gap, i.e., $\hat{D}^{\star}<P^{\star}$, under certain conditions this gap can be bounded.
@@ 4.1 Algorithm
  Return: $\boldsymbol{\theta}$, $\boldsymbol{\lambda}$.

### v2

@@ 1 Introduction
Predicting multiple future values of time series data, also known as multi-step forecasting (Bontempi et al., 2013), has a myriad of applications such as forecasting weather (Wang et al., 2016), electricity demand (Yi et al., 2022), prices (Chen et al., 2018), and passenger demand  (Bai et al., 2019). Several approaches to generating predictions for the next window have been proposed, including direct (Chevillon, 2007; Sorjamaa et al., 2007), autoregressive or recursive (Hill et al., 1996; Tiao & Tsay, 1994; Hamzaebi et al., 2009) and MIMO techniques (Kline, 2004; Bontempi, 2008; Bontempi & Ben Taieb, 2011; Kitaev et al., 2019; Zhou et al., 2021; Wu et al., 2021; Liu et al., 2021; Zhou et al., 2022b; Nie et al., 2022; Zeng et al., 2023; Garza & Mergenthaler-Canseco, 2023; Das et al., 2023b; Liu et al., 2024, 2022). Moreover, a plethora of learning parametrizations exist, ranging from linear models (Box & Jenkins, 1976; Zeng et al., 2023) to recent transformer  (Kitaev et al., 2019; Zhou et al., 2021; Wu et al., 2021; Liu et al., 2021, 2024, 2022) and custom (Zhou et al., 2022a) architectures.
@@
Recent works using transformer based architectures focus on aggregate metrics, see for example (Kitaev et al., 2019; Wu et al., 2021; Nie et al., 2022), while addressing errors at different time steps has recieved little attention (Cheng et al., 2023).
However, having no control on the error incurred at each time step in the predicted window can be detrimental in scenarios that can be affected by this variability. For example, analysing average behaviour is insufficient to assess financial risks in econometrics  (Chavleishvili & Manganelli, 2019) or to ensure stability when using the predictor in a Model Predictive Control framework (Terzi et al., 2021).
Therefore, our first contribution is addressing this problem through constrained learning.
@@
Since this constrains the distribution of the loss per time step across the window, we call this approach loss shaping constraints. Due to the challenges in finding appropriate constraints, we leverage a recent approach (Hounie et al., 2023) to re-interpret the specification as a soft constraint and find a minimal norm relaxation while jointly solving the learning task. We leverage recent duality results (Chamon & Ribeiro, 2020) to show that, despite being non-convex, the resulting problem has a bounded duality gap. This allows us to provide approximation guarantees at the step level.
@@ 1.1 Related work
Although our method is related to other loss shaping approaches (Park et al., 2023; Wang et al., 2024; Cheng et al., 2023) in the sense that it alters the weighting given to each time step on the loss, both the motivation and the problems that these approaches address differ from our work. (Park et al., 2023) aim to downweight the impact of large errors as long as they have low probability, regardless of their position in the predictive window. Our approach aims to control average errors at each step and can thus go in the opposite direction of sacrificing performance in order to reduce undesirably large errors at specific steps.  (Wang et al., 2024) gives each time-step a fixed weighting that is inversely proportional to their position in the window, thus aiming to discount the impact of far-future errors and give a higher weighting to closer steps. However, imposing a fixed weighting does not take into account the difficulty of fitting an error at a certain time-step, which will depend also on the model and the data. That is, imposing a certain weighting across the window need not result in the desired distribution of errors.
Lastly, recent works in generative time series models have also sought to impose constraints using penalty based methods (Coletta et al., 2023), but the nature of the constraints and proposed approach also differs from ours.
@@ 2 Multi-Step Time Series Forecasting
$~{}\min_{\boldsymbol{\theta}\in\Theta}~{}\mathbb{E}_{(\mathbf{x}_{c},y_{p})\sim\mathfrak{D}}\left[\frac{1}{T_{p}}\sum_{i=1}^{T_{p}}\ell_{i}({f_{\boldsymbol{\theta}}(\mathbf{x}_{c})},y_{p})\right],$ (ERM)
@@
However, this choice of objective does not account for the structure or distribution of the errors across different time steps, which can lead to disparate behaviour across the predicted window as depicted in Figure 1. In particular, we observe empirically that state-of-the-art transformer architectures (Kitaev et al., 2019; Zhou et al., 2021; Wu et al., 2021; Liu et al., 2021; Zhou et al., 2022a) can yield highly varying loss dynamics, including non-monotonic, flat and highly non-linear landscapes, as presented in Appendix D inline][IH]: Add old shape of ERM section to appendix
@@ 3 Loss Shaping Constraints
3.1 Adapting Constraints: Resilient Constrained Learning
In the unconstrained risk minimization problem  (ERM) an optimal function $\theta^{\star}$ always exists. This is not the case of the constrained learning problem in (3) in which there may be no parameters in $\Theta$ that satisfy all the requirements if the target losses $\epsilon_{i}$ are too restrictive.
In practice, landing on satisfiable loss shaping requirements may necessitate the relaxation of some constraints –i.e., the values of $\boldsymbol{\epsilon}_{i}$ in (3)–. This is challenging because assessing the impact of tightening or relaxing a particular constraint can have intricate dependencies with the model class, the unknown data distribution, and learning algorithm, which are hard to determine a priori.
Thus, the point at issue is coming up with reasonable constraints that achieve a desirable trade-off between controlling the distribution of the error across the predictive window and attaining good average performance. That is, for larger constraint levels the average performance improves, although the constraints will have less impact on the optimal function $f_{\theta^{\star}}$.
In order to do so, we introduce a non negative perturbation $\boldsymbol{\zeta}_{t}\in\mathbb{R}_{+}$ associated to each time step, and consider relaxing the $i-th$ constraint in the original problem by $\boldsymbol{\zeta}_{i}$. Explicitly, we impose $\mathbb{E}_{(\mathbf{x}_{c},y_{p})\sim\mathfrak{D}}\left[\ell_{i}({f_{\boldsymbol{\theta}}(\mathbf{x}_{c})},y_{p})\right]\leq\boldsymbol{\epsilon}_{i}+\boldsymbol{\zeta}_{i}$. We also introduce a differentiable, convex, non-decreasing cost $h:\mathbb{R}_{+}^{T_{p}}\to\mathbb{R}_{+}$, that penalizes deviating from the original specification, e.g., the squared $L_{2}$ norm $h(\boldsymbol{\zeta})\propto\|\boldsymbol{\zeta}\|_{2}^{2}$.Thus, we seek a relaxation $\boldsymbol{\zeta}^{\star}$ that achieves a desirable trade-off by equating marginal decrease in the objective and the marginal increase in relaxation cost. Explicitly,
$~{}-\partial h(\boldsymbol{\zeta}^{\star})\in\partial P^{\star}(\boldsymbol{\epsilon}+\boldsymbol{\zeta}^{\star}),$ (1)
where $\partial P^{\star}$ denotes the subdifferential of $P^{\star}$.
This allows us to re-interpret the initial constraint levels $\boldsymbol{\epsilon}$ as a soft constraint, and learn a relaxation that makes the problem easier to solve. As shown by (Hounie et al., 2023) this relaxation can be found while jointly solving the learning task by solving the problem
$\min_{\boldsymbol{\zeta}\in\mathbb{R}^{T_{p}}}P^{\star}(\boldsymbol{\epsilon}+\boldsymbol{\zeta})+h(\boldsymbol{\zeta})\;.$ (R-LS)
Further discussion on equivalent formulations that provide a straightforward way to compute the resilient relaxation are discussed in Appendix B.2. In the next section, we introduce practical primal-dual algorithms that enable approximating the constrained (3) and resilient (B.2) problems based on samples and finite (possibly non-convex) parametrizations such as transformer architectures.
@@ 4 Empirical Dual Resilient and Constrained Learning
A challenge in solving (3) and (B.2) is that in general, (i) no closed form projection into the feasible set or proximal operator exists, and (ii) it involves an unknown data distribution $\mathfrak{D}$. In what follows, we describe our approach for tackling the Resilient problem, noting that Primal-Dual algorithms tackling the original constrained problem can be similarly derived by simply excluding the slack variables.
To undertake B.2, (i) we replace expectations by sample means over a data set $\{(\boldsymbol{x}_{c}^{n},y_{p}^{n})\>:\>n=1,\cdots,N\}$, as typically done in (unconstrained) statistical learning, and (ii) resort to its Lagrangian dual.
@@
where $\hat{L}$ is the empirical Lagrangian of (B.2), defined as
${\hfilneg\hat{L}(\boldsymbol{\boldsymbol{\theta}},\boldsymbol{\lambda},\boldsymbol{\epsilon},\boldsymbol{\zeta}):=\frac{1}{2}h(\boldsymbol{\zeta})+\hskip 10000.0pt}$
@@
The Empirical Dual problem (ED-LS) is an approximation of the Dual problem associated to B.2 based on training samples. The dual problem itself can be interpreted as finding the tightest lower bound on the primal. Although the estimation of expectations using sample means and non-convexity of the hypothesis class can introduce a duality gap, i.e., $\hat{D}^{\star}<P^{\star}$, under certain conditions this gap can be bounded.
@@ 4.1 Algorithm
  Return: $\boldsymbol{\theta}$, $\boldsymbol{\lambda}$, $\boldsymbol{\zeta}$.
Figure 2: Test MSE at each prediction step for two datasets across different models, datasets and predictive windows. The top row shows results for the Weather dataset with a predictive window length of 96 steps, and the second row corresponds to the Exchange Rate dataset with a predictive window length of 720 steps. Each column corresponds to a different architecture, and each curve represents a different training algorithm, we include both the ERM baseline and our method using a constant constraint across the window for all models.